OpenStudy (anonymous):

use the Maclaurin series to write a symbolic representation for e^x including the general form and the first six terms

OpenStudy (anonymous):

\[1+x+\frac{x^2}{2}+\frac{x^3}{3!}+\frac{x^4}{4!}+\frac{x^5}{5!}+\frac{x^6}{6!}\]

OpenStudy (anonymous):

thanks